Question 6

According to a recent survey, 17% of the population is left handed. If you randomly select 6 people what is the probability that at least one of them is left handed?

Round your answer to three decimal places.

Question 71 pts

If  LaTeX: P(A) = 0.45P ( A ) = 0.45  then for any event  LaTeX: BB we have  LaTeX: P(A\cap B) \leq 0.45P ( A ∩ B ) ≤ 0.45

Group of answer choices

True

False
Answer #1

Solution:

6. Given that p = 0.17, n = 6

P(X >= 1) = P(1) + P(2) + P(3) + P(4) + P(5) + P(6)
= 0.4018 + 0.2057 + 0.0562 + 0.0086 + 0.0007 + 0.0000
= 0.6730

Binomial distribution: P(X = r) = nCr*p^r*q^(n-r)
Ex = P(X = 1) = 6C1*0.17^1*0.83^5
= 0.4018

7. option B. false.